OpenStudy (anonymous):

You want to buy a new car in 8 years and expect the car to cost 97,000. Your bank offers a plan with a guaranteed APR of 4.5%. If you make regular monthly deposits each month. How much should you deposit each month to end up with 97,000 in 8 years? What formula do I use?

OpenStudy (anonymous):

Do I use A= p(1+APR/n)^(ny)

OpenStudy (wolf1728):

Total = Principal * (1 + Rate) ^ Years

OpenStudy (wolf1728):

I'm guessing n means number of months.

OpenStudy (wolf1728):

Hmmm, I'm beginning to think this is more of an annuity type investment instead of a compound interest problem.

OpenStudy (anonymous):

i think i use the PMT formula

OpenStudy (wolf1728):

I think I finally have it:

OpenStudy (wolf1728):

No that is the formula to use: the amount is the amount deposited every month Hmmm, I guess I'll have to solve that for amount.

OpenStudy (anonymous):

I got 841.30 tell me if u got that

OpenStudy (wolf1728):

As of the moment I'm trying to solve that formula foramount

OpenStudy (anonymous):

kk let me know when u have an answer

OpenStudy (anonymous):

Did u solve it

OpenStudy (amistre64):

i dont use formulas, i use concepts to develop my own formulas ....

OpenStudy (amistre64):

You want to buy a new car in 8 years and expect the car to cost 97,000. Your bank offers a plan with a guaranteed APR of 4.5%. If you make regular monthly deposits. How much should you deposit each month to end up with 97,000 in 8 years 8*12 payments = 96 total deposits spose we had to take out the loan now and pay it back in 8 years? \[B_o = B_o\] \[B_1 = B_ok-P\] \[B_2 = B_ok^2-Pk-P\] \[B_3 = B_ok^3-Pk^2-Pk-P\] ...\[B_n=B_ok^n-P\frac{1-k^n}{1-k}\] solving for P, when B_n = 0 \[B_ok^n=P\frac{1-k^n}{1-k}\] \[B_ok^n\frac{1-k}{1-k^n}=P\]

OpenStudy (amistre64):

P = 97000k^(96)(1-k)/(1-k^(96)), k=1+.045/12 you need to make deposits of abt 1205.05 each month to save up that amount of cash in 8 years

OpenStudy (amistre64):

or, if i try this another way to verify my idea ...

OpenStudy (amistre64):

\[B_0 = P\] \[B_1 = Pk+P\] \[B_2 = Pk^2+Pk+P\] \[B_3 = Pk^3+Pk^2+Pk+P\] ...\[B_n = Pk+P\frac{1-k^n}{1-k}\]

OpenStudy (amistre64):

\[97000=Pk^{96}+P\frac{1-k^{96}}{1-k}\] \[97000=P(k^{96}+\frac{1-k^{96}}{1-k})\] \[97000\div (k^{96}+\frac{1-k^{96}}{1-k})=P\]

OpenStudy (amistre64):

thats better, lets try 252.35 a month

OpenStudy (amistre64):

getting closer :) had a typo in the calculator

OpenStudy (amistre64):

thats does seem to work out rather nicely